OpenStudy (anonymous):

Four times the sum of a number and 3 is at least 28

OpenStudy (shamil98):

4(x+3) >= 28

OpenStudy (anonymous):

steps: 4x+12>=28 4x>=28-12 4x>=16 x>=4

OpenStudy (anonymous):

Umm I think u have to take parenthesis offhand subtract 3 on both sides then 28 - 3 is 25 then idk after that

OpenStudy (anonymous):

Oh thanks